OpenStudy (kaythxbai):

margo boils 500 ml of water to make 80 g of the pasta. if margo boils 2500 ml of water, how many grams of pasta can she make?

jhonyy9 (jhonyy9):

use the rule of 3 simple and so you can writing 500 ml of water to make 80 g pasta 2500 ml -------------- x -------------------------------- x = 2500*80/500 = ? g

jhonyy9 (jhonyy9):

simplify by 100 and will get x = 25*80/5 simplify by 5 and will get x = 5*80 = ? g

jhonyy9 (jhonyy9):

ok. ?

jhonyy9 (jhonyy9):

so hope helped was my pleasure good luck bye bye